You get a box of assorted chocolates for Christmas. There are the following types; 8 dark chocolate-coated caramels, 7 milks chocolate-coated caramels, 7 dark chocolate-coated nuts, 6 milk chocolate-coated vanilla creams, 8 dark chocolate-coated cherries, 7 milk chocolate-coated strawberry creams. Determine the following probabilities;
a) of not getting a milk chocolate-covered piece
b) of getting caramel
c) of getting a milk chocolate-covered piece or a caramel
d) of getting a nut or a caramel
e) of getting 2 dark chocolate-coated cherries in a row and then 5
caramels in a row.

Answer by Ultimatemagic(23)

A= 20/43 B= 15/43 C= 35/43 D=22/43